How many total atoms are in .170g of P2O5?How many total atoms are in .170g of P2O5?
Answer #1

No. of atoms in one mole of atoms= 6.022×1023

weight of one mole of P2O5= 283.89 g

no.of moles of P2O5=0.17/283.89

= 5.98 x 10^-4

no. of atoms= 5.98 x 10^-4 *  6.022×1023 = 36.011 x 10^18 = 3.6011 x 10^19 atoms